OpenStudy (anonymous):

solve each of the following in quadratic equations by factoring: x squared + 2x 2x squared = 6x 2a squared + a =0 6m squared + 9m=0 t(2t-3) = 2 what will be the answer

OpenStudy (anonymous):

what is the question ?

OpenStudy (anonymous):

quadratic equations by factoring ..........

OpenStudy (anonymous):

so is it like x(x+2)... ???

OpenStudy (anonymous):

yes .. so what are the answer .. ?

OpenStudy (anonymous):

1st, x(x+2), 2nd, 2x(x-3), 3rd, a(2a+1), 4th, 3m(2m+3), and final (2t+1)(t-2)

OpenStudy (anonymous):

i have more !

OpenStudy (anonymous):

can u answer more .. ?

OpenStudy (anonymous):

unfortunately :(

OpenStudy (anonymous):

why .. ?

OpenStudy (diyadiya):

@burakb Please show him/her how to do the question, rather than giving just the answers :)

OpenStudy (anonymous):

ok .

OpenStudy (anonymous):

correct !

OpenStudy (diyadiya):

@lilyi did you understand how he got the answers?

OpenStudy (anonymous):

uhhhmmm .. i dont know ?

OpenStudy (anonymous):

.. uhhhmm ... how was it ?

OpenStudy (diyadiya):

x squared + 2x \[x^2+2x\] now x is common in both terms, so factor out x x(x+ 2)

OpenStudy (anonymous):

ok .

OpenStudy (anonymous):

how about the step 2 .. ?

OpenStudy (diyadiya):

there's only one step :) that's the answer

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

but the answer are not to be improve .. !

OpenStudy (diyadiya):

Hmm what? improve?

OpenStudy (anonymous):

because thats not a exactly the answer !

OpenStudy (diyadiya):

then, what is the answer given in your book?

OpenStudy (anonymous):

read burakb answer !

OpenStudy (diyadiya):

"1st, x(x+2) " this is what he posted and that's the same thing i got! :D

OpenStudy (anonymous):

yup .. buuuuuuutttttttt ..........

OpenStudy (diyadiya):

but.. ? Well, if you have more questions please do post it :)

OpenStudy (anonymous):

im not right .. , what will be the solution of that !

OpenStudy (anonymous):

thats i mean so !

OpenStudy (diyadiya):

Sorry, i didn't get you! What is your question now?

OpenStudy (anonymous):

a(2a+1) .., what are the solution .. ?

OpenStudy (diyadiya):

its already factored

OpenStudy (callisto):

''solve each of the following in quadratic equations by factoring'' => factor => solve For example, in question one, \[x^2 + 2x=0\]Factor the left side by taking our the common factor x. Then you get: x(x+2) = 0 To solve the equation, put each factor = 0. Since if one of the factor = 0, 0 times other factor can give 0 as the answer, that is the right side. So, x = 0 or x+2 = 0 <- solve these two equations, you'll get x = 0 or x = -2 Got it?

OpenStudy (anonymous):

tnx .. yes !

OpenStudy (anonymous):

what are the answer .. ?

OpenStudy (callisto):

Can you figure it out?

OpenStudy (anonymous):

bbye ..!